# publications.bib

### v3

@@ 2.1.2. Compactifications
To each compatible choice $\Sigma=\{\Sigma_{\Phi}\}$ of admissible smooth rational polyhedral cone decomposition data, there is an associated proper smooth algebraic scheme $\mathfrak{S}(G,X)_{K}^{\mathrm{tor}}$ over $\mathcal{O}_{E,\mathfrak{p}}$ containing $\mathfrak{S}(G,X)_{K}$ as an open dense subscheme, together with a semiabelian family $\mathcal{A}$ over $\mathfrak{S}(G,X)_{K}^{\mathrm{tor}}$. Moreover, we have a stratification
$\mathfrak{S}(G,X)_{K}^{\mathrm{tor}}=\coprod_{(\Phi,\sigma)}Z_{(\Phi,\sigma)},$
where $\sigma$ is a face of $P_{\Phi}$.
Finally, observe that there is a cusp label corresponding to taking the filtration $Z_{-2}=0,Z_{-1}=L\otimes R$, and the torus argument $X=Y=0$. This trivial cusp label will correspond to the original Shimura variety $\mathfrak{S}(G,X)_{K}$ in the stratification above.
@@ 2.2. Igusa Varieties
We fix now some $b\in B(G,\mu)$, and consider a geometric point $x\in\mathfrak{S}(\mathbf{G},X)_{K}(\overline{\mathbb{F}}_{p})$ lying in the Newton strata for $b$. This corresponds to some abelian variety $\mathcal{A}_{x}$, which has $p$-divisible group with $G$-structure $\mathbb{X}:=\mathcal{A}_{x}[p^{\infty}]$ given by $b$. Up to replacing $x$ by another element in its isogeny class, we can assume $\mathbb{X}$ is completely slope divisible. Thus, we can write $\mathbb{X}=\oplus_{i=1}^{r}\mathbb{X}_{i}$, where the $\mathbb{X}_{i}$ are isoclinic $p$-divisible groups of strictly decreasing slopes.
@@
[…]
where we denote by $\mathfrak{S}(\mathbf{G},X)_{K,\overline{\mathbb{F}}_{p}}$ the (geometric) special fiber of $\mathfrak{S}(\mathbf{G},X)_{K}$. This turns out to be a closed subset of $\mathfrak{S}(\mathbf{G},X)_{K,\overline{\mathbb{F}}_{p}}$, and thus we can give this subset the induced reduced scheme structure. We will continue to denote the associated scheme by $\mathscr{C}_{\mathbb{X}}$, and it turns out this scheme is smooth.
Let $\mathscr{G}$ be the $p$-divisible group of the restriction to $\mathscr{C}_{\mathbb{X}}$ of the universal abelian variety over $\mathfrak{S}(\mathbf{G},X)_{K}$. We further define $\mathrm{Ig}^{b}$ as the scheme over $\mathscr{C}_{\mathbb{X}}$ parametrizing, for any perfect $\mathscr{C}_{\mathbb{X}}$-scheme $\mathscr{T}$, isomorphisms $\mathscr{G}\times_{\mathscr{C}_{\mathbb{X}}}\mathscr{T}\simeq\mathbb{X}\times_{\overline{\mathbb{F}}_{p}}\mathscr{T}$ which preserve $G$-structure. Equivalently, we can define $\mathrm{Ig}^{b}$ as the functor sending an $\overline{\mathbb{F}}_{p}$-algebra $R$ to the set
[…]
By [CS1, Corollary 4.3.5], we know that this scheme is perfect, and hence it lifts uniquely to a flat $p$-adic formal scheme, which we denote by $\mathfrak{Ig}^{b}$ over $\mathrm{Spf}(W(\overline{\mathbb{F}}_{p}))$.
We write $\mathfrak{Ig}_{C}^{b}$ for the perfectoid space attached to the adic generic fiber of $\mathfrak{Ig}^{b}$ over $C$. These spaces are supposed to model the fibers of the Hodge-Tate period morphism, a connection we will elaborate upon in §3.
@@ 2.2.1. Compactifications
In order to understand (partial) minimal and toroidal compactifications of $\mathrm{Ig}^{b}$, we must first consider compactifications of the central leaf $\mathscr{C}_{\mathbb{X}}$. As in the discussion in [CS2, §3.1], the central leaf $\mathscr{C}_{\mathbb{X}}$ is a well-positioned subset of $\mathfrak{S}(\mathbf{G},X)_{K,\overline{\mathbb{F}}_{p}}$, and thus admits partial toroidal and minimal compactifications, which we will denote by $\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}$ and $\mathscr{C}_{\mathbb{X}}^{*}$ respectively. Moreover, let $Z$ be a cusp label at level $K(N)$. This determines a locally closed boundary stratum $\mathscr{C}_{\mathbb{X},Z}\subset\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}$.
@@
The Igusa variety $\mathrm{Ig}^{b}$ over $\mathscr{C}_{\mathbb{X}}$ extends to a perfect scheme $\mathrm{Ig}^{b,\mathrm{tor}}$ over $\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}$. More precisely, we can define $\mathrm{Ig}^{b,\mathrm{tor}}$ as follows. Let $\mathcal{A}$ denote the universal semi-abelian scheme over $\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}$. This is the restriction to $\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}$ of the universal semi-abelian scheme over $\mathfrak{S}(\mathbf{G},X)_{K,\overline{\mathbb{F}}_{p}}$. Then, we know from [CS2, Proposition 3.2.1] that the connected part $\mathcal{A}[p^{\infty}]^{\circ}$ of $\mathcal{A}[p^{\infty}]$ is a $p$-divisible group. Moreover, if we denote by $\mathcal{A}[p^{\infty}]^{\mu}$ the multiplicative part, then this is also a $p$-divisible group. We thus let $\mathcal{A}[p^{\infty}]^{(0,1)}=\mathcal{A}[p^{\infty}]^{\circ}/\mathcal{A}[p^{\infty}]^{\mu}$ be the biconnected part. We can similarly define $\mathbb{X}^{\circ},\mathbb{X}^{(0,1)}$ as the connnted and biconnected parts of $\mathbb{X}$.
@@
Thus, we can define $\mathrm{Ig}^{b,\mathrm{tor}}$ to be the scheme which, for a perfect $\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}$-scheme $\mathscr{T}$, parametrizes $\mathcal{O}_{B}$linear isomorphisms $\rho:\mathcal{A}[p^{\infty}]^{\circ}\times_{\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}}\mathscr{T}\xrightarrow{\sim}\mathbb{X}^{\circ}_{\mathscr{T}}$ and a scalar in $\mathbb{Z}^{\times}_{p}(\mathscr{T})$ such that the induced isomorphism $\rho^{(0,1)}:\mathcal{A}[p^{\infty}]^{(0,1)}\times_{\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}}\mathscr{T}\xrightarrow{\sim}\mathbb{X}_{\mathscr{T}}^{(0,1)}$ obtained by quotienting by the multiplicative parts commutes with the polarizations up to the given element of $\mathbb{Z}_{p}^{\times}(\mathscr{T})$. Here, $\mathbb{Z}_{p}^{\times}(\mathscr{T})$ is the set of $\mathscr{T}$-points of the group scheme $\mathbb{Z}_{p}^{\times}\times\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}$.
@@
Finally, we define the partial minimal compactification $\mathrm{Ig}^{b,*}$ as the normalization of $\mathscr{C}_{\mathbb{X}}^{*}$ in $\mathrm{Ig}^{b}$. Since we have $\mathrm{Ig}^{b}\subset\mathrm{Ig}^{b,*},\mathrm{Ig}^{b,\mathrm{tor}}$, we will denote the boundaries by $\partial\mathrm{Ig}^{b,*}$ and $\partial\mathrm{Ig}^{b,\mathrm{tor}}$, respectively. These schemes are all perfect, and we can lift them to $p$-adic formal schemes over $\mathrm{Spf}(W(\overline{\mathbb{F}}_{p}))$. We similarly denote by $\mathfrak{Ig}_{C}^{b,*}$, $\mathfrak{Ig}_{C}^{b,\mathrm{tor}}$, $\partial\mathfrak{Ig}_{C}^{b,*}$, and $\partial\mathfrak{Ig}_{C}^{b,\mathrm{tor}}$ the associated perfectoid spaces over $C$.

### v5

@@ 2.1.3. Compactifications
To each compatible choice $\Sigma=\{\Sigma_{\Phi}\}$ of admissible smooth rational polyhedral cone decomposition data, there is an associated flat proper normal algebraic scheme $\mathscr{S}(G,X)_{K}^{\mathrm{tor}}$ over $\mathcal{O}_{E,\mathfrak{p}}$ (which is even smooth when $p\nmid N$) containing $\mathscr{S}(G,X)_{K}$ as an open dense subscheme, together with a semiabelian family $\mathcal{A}$ over $\mathscr{S}(G,X)_{K}^{\mathrm{tor}}$. Moreover, we have the following
  1. (1)
    We have a decomposition into reduced locally closed subvarieties
    $\mathscr{S}(\mathbf{G},X)_{K}^{\mathrm{tor}}=\bigsqcup_{[Z]}\mathscr{S}(\mathbf{G},X)^{\mathrm{tor}}_{K,Z},$
    where $Z=(Z,\Phi)$ is a cusp label of level $K$, and where each $\mathscr{S}(\mathbf{G},X)^{\mathrm{tor}}_{K,Z}$ is flat over $O_{E_{\mathfrak{p}}}$.
  2. (2)
    There exists a scheme
    $C_{Z}\rightarrow\mathscr{S}_{Z}$
    (which when $p\nmid N$ is an abelian scheme torsor). Moreover, there exists a torsor under the torus $E_{\Phi}$, $\Xi_{Z}\rightarrow C_{Z}$, and a relative toroidal embedding $\Xi_{Z}\hookrightarrow\Xi_{Z,\Sigma_{Z}}$ induced by the cone $\Sigma_{Z}$ (see [41, Definition 6.1.2.3] for the definition of relative toroidal embedding), as depicted in the following diagram:
    $\Xi_{Z}$$C_{Z}$$\mathscr{S}_{Z}$$\Xi_{Z,\Sigma_{Z}}$}
    and $\Xi_{Z}$ has an action of $\Gamma_{Z}$ which extends to $\Xi_{Z,\Sigma_{Z}}$. We denote $\Xi_{Z,\Sigma_{Z}}\backslash\Xi_{Z}$ by $\partial\Xi_{Z,\Sigma_{Z}}$.
    If we let
    $\left(\mathscr{S}(\mathbf{G},X)^{\mathrm{tor}}_{K}\right)^{\wedge}_{[Z]},$
    denote the completion of $\mathscr{S}(\mathbf{G},X)_{K}^{\mathrm{tor}}$ along the locally closed component indexed by [Z][Z] [^2] By convention, by the formal completion of a scheme $X$ along a locally closed subscheme $Z$ we mean the formal completion of the open subscheme $X-(\overline{Z}-Z)$ along its closed subscheme $Z$., then we have an isomorphism
    $\left(\mathscr{S}(\mathbf{G},X)^{\mathrm{tor}}_{K}\right)^{\wedge}_{[Z]}\simeq\mathfrak{X}_{Z,\Sigma_{Z}}/\Gamma_{Z},$
    where $\mathfrak{X}_{Z,\Sigma_{Z}}$ is the completion of $\Xi_{Z,\Sigma_{Z}}$ along $\partial\Xi_{Z,\Sigma_{Z}}$.
We note that $C_{Z},\mathscr{S}_{Z},\Xi_{Z}$ parameterize certain degeneration structures on semiabelian schemes, which we will briefly sketch.
We first recall some facts about degenerations of abelian schemes, from [41, §3.3] and [8, §2.5.1]. Let $R$ be a normal local ring with fraction field $H$. Consider a polarized abelian variety $(A,\lambda)$ over $H$ with $\mathcal{O}_{B}$-structure, and $\mathcal{A}$ a semiabelian scheme over $R$ which is a degeneration of $A$ (that is, $\mathcal{A}_{H}\simeq A$). Then this uniquely determines a short exact sequence
where $T$ is a torus, $\mathcal{B}$ is an abelian scheme over $R$, and $\mathcal{G}$ is the Raynaud extension (see [41, Definition 3.3.3.9] for definition). Let $X=\mathbb{X}^{*}(T)$, this is a free module over $R$. The lattice $X$ has an action of $\mathcal{O}_{B}$, and so does $\mathcal{B}$. Then, $\mathcal{G}$ determines, and is uniquely determined by, an $\mathcal{O}_{B}$-linear map $c:X\rightarrow\mathcal{B}^{\vee}$. Similarly, we can consider a degeneration $\mathcal{A}^{\vee}$ over $R$ of the dual $A^{\vee}/H$, which gives us a short exact sequence
and if we similarly let $Y=\mathbb{X}^{*}(T^{\vee})=\mathbb{X}_{*}(T)$, the extension $\mathcal{G}$ determines, and is uniquely determined by an $\mathcal{O}_{B}$-linear map $c^{\vee}:Y\rightarrow\mathcal{B}$.
Moreover, note that, as shown in [41, §3.4], the polarization on $A$ extends uniquely to a homomorphism $\lambda:\mathcal{G}\rightarrow\mathcal{G}^{\vee}$. The homomorphism $\lambda$ determines and is uniquely determined by the data of the polarization on $\mathcal{B}$, and an injective $\mathcal{O}_{B}$-linear map $\phi:Y\rightarrow X$.
Then, given a cusp label $Z$ of level $K(N)$, we can define a smaller-dimensional Shimura variety $\mathscr{S}_{Z}$ which will be a union of connected components of a PEL moduli space which parameterizes the data of abelian schemes $\mathcal{B}$ as described above, over which we consider another moduli space $C_{Z}$, additionally parametrizing the maps $c,c^{\vee}$, as well as extensions of maps $f_{0}:\frac{1}{N}X\rightarrow\mathcal{B}^{\vee}$ and $f_{0}^{\vee}:\frac{1}{N}Y\rightarrow\mathcal{B}$. The space $\Xi_{Z}$ parameterizes liftings of the map $f^{\vee}:\frac{1}{N}Y\rightarrow\mathcal{B}$ to a map $f^{\vee}:\frac{1}{N}Y\rightarrow\mathcal{G}$.
Finally, we also want to consider the integral model over $\mathcal{O}_{E_{\mathfrak{p}}}$ of the minimal compactification which was constructed in [41, §7.2.3], and we denote this by $\mathscr{S}(\mathbf{G},X)_{K}^{*}$. We will denote the generic fibers of the toroidal and minimal compactifications by $\mathrm{Sh}(\mathbf{G},X)_{K}^{\mathrm{tor}}$ and $\mathrm{Sh}(\mathbf{G},X)^{*}_{K}$, respectively.
@@ 2.2. Igusa Varieties
We now turn our attention to studying Igusa varieties. Throughout this subsection the level $K=K(N)$ satisfies $p\nmid N$. We start with the basic definition.
2.2.1. The Basic Definition
Definition 2.6.
A $p$-divisible group $\mathbb{X}$ with $\mathbf{G}_{\mathbb{Z}_{p}}$-structure is a tuple $(\mathbb{X},\lambda,\iota)$ where $\mathbb{X}$ is a $p$-divisible group of dimension $n[F:\mathbb{Q}]$, $\lambda$ is a quasi-polarization (i.e. a quasi-isogeny $\lambda:\mathbb{X}\dashrightarrow\mathbb{X}^{\vee}$), and $\iota:\mathcal{O}_{B_{\mathbb{Q}_{p}}}\hookrightarrow\mathrm{End(\mathbb{X})}$, where $\mathcal{O}_{B_{\mathbb{Q}_{p}}}$ is the $p$-adic completion of $\mathcal{O}_{B}$ in $B_{\mathbb{Q}_{p}}$, such that $\lambda\circ\iota(b^{*})=\iota(b)^{\vee}\circ\lambda$.
For notational simplicity, we will often drop $\lambda,\iota$ from the notation.
Let $\mu\in\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})^{+}$ be the dominant cocharacter induced by the inverse of the Hodge cocharacter associated with the conjugacy class $X$. We fix now some $b\in B(G,\mu)$, and consider a geometric point $x\in\mathscr{S}(\mathbf{G},X)_{K}(\overline{\mathbb{F}}_{p})$ lying in the Newton strata for $b$. This defines an abelian variety $\mathcal{A}_{x}/\mathop{\rm Spec}(\overline{\mathbb{F}}_{p})$. Consider now the associated $p$-divisible group $\mathbb{X}:=\mathcal{A}_{x}[p^{\infty}]$, the $p$-divisible group $\mathbb{X}$ has $\mathbf{G}_{\mathbb{Z}_{p}}$-structure acquired from the additional structures on $\mathcal{A}_{x}$.
Up to replacing $x$ by another element in its isogeny class, we can assume $\mathbb{X}$ is completely slope divisible, following a result of Oort and Zink [49, Theorem 2.1]. Thus, we can write $\mathbb{X}=\oplus_{i=1}^{r}\mathbb{X}_{i}$, where the $\mathbb{X}_{i}$ are isoclinic $p$-divisible groups of strictly decreasing slopes.
@@ 2.2.1. The Basic Definition
$\mathscr{C}_{\mathbb{X}}:=\{x\in\mathscr{S}(\mathbf{G},X)_{K,\overline{\mathbb{F}}_{p}}:\exists\text{ isomorphism }\rho:\mathcal{A}_{x}[p^{\infty}]\times k(\overline{x})\simeq\mathbb{X}\times k(\overline{x})\text{ preserving $\mathbf{G}_{\mathbb{Z}_{p}}$-structure}\},$
known as the central leaf, where we denote by $\mathscr{S}(\mathbf{G},X)_{K,\overline{\mathbb{F}}_{p}}$ the (geometric) special fiber of $\mathscr{S}(\mathbf{G},X)_{K}$. It is shown in [46, Proposition 1] that this is a locally closed subset of $\mathscr{S}(\mathbf{G},X)_{K,\overline{\mathbb{F}}_{p}}$, and that we can give this subset the induced reduced scheme structure which makes the associated scheme $\mathscr{C}_{\mathbb{X}}$ smooth.
Let $\mathscr{G}$ be the $p$-divisible group of the restriction to $\mathscr{C}_{\mathbb{X}}$ of the universal abelian variety over $\mathscr{S}(\mathbf{G},X)_{K}$. We further define $\mathrm{Ig}^{b}$ to be the functor parametrizing, for any perfect $\mathscr{C}_{\mathbb{X}}$-scheme $\mathscr{T}$, isomorphisms $\mathscr{G}\times_{\mathscr{C}_{\mathbb{X}}}\mathscr{T}\simeq\mathbb{X}\times_{\overline{\mathbb{F}}_{p}}\mathscr{T}$ which preserve $\mathbf{G}_{\mathbb{Z}_{p}}$-structure. Equivalently, we can define $\mathrm{Ig}^{b}$ as the functor sending an $\overline{\mathbb{F}}_{p}$-algebra $R$ to the set
(3) $\mathrm{Ig}^{b}(R)=\{(\rho,x):x\in\mathscr{S}(\mathbf{G},X)_{K}(R),\rho:\mathcal{A}_{x}[p^{\infty}]\xrightarrow{\sim}\mathbb{X}_{R}\text{ preserving $\mathbf{G}_{\mathbb{Z}_{p}}$-structure}\}.$
By [7, Proposition 4.3.3], we know that this functor is representable, and by [7, Corollary 4.3.5] the scheme is perfect, and hence it lifts uniquely to a flat $p$-adic formal scheme, which we denote by $\mathfrak{Ig}^{b}$ over $\mathrm{Spf}(W(\overline{\mathbb{F}}_{p}))$. Moreover, from [8, Corollary 2.3.2], we know that $\mathrm{Ig}^{b}\rightarrow\mathscr{C}_{\mathbb{X}}$ is a torsor for $\mathrm{Aut}_{G}(\mathbb{X})_{\overline{\mathbb{F}}_{p}}$, the formal group scheme of automorphisms of $\mathbb{X}$ compatible with extra structures viewed as an fpqc sheaf over $\overline{\mathbb{F}}_{p}$.
We write $\mathfrak{Ig}_{C}^{b}$ for the perfectoid space attached to the adic generic fiber of $\mathfrak{Ig}^{b}$ over $C$. These spaces are supposed to model the fibers of the Hodge-Tate period morphism, a connection we will elaborate upon in §3.1.
@@ 2.2.2. Compactifications
In order to understand (partial) minimal and toroidal compactifications of $\mathrm{Ig}^{b}$, we must first consider compactifications of the central leaf $\mathscr{C}_{\mathbb{X}}$. As constructed [54, §3.2.1], the central leaf $\mathscr{C}_{\mathbb{X}}$ admits partial toroidal and minimal compactifications, which we will denote by $\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}$ and $\mathscr{C}_{\mathbb{X}}^{*}$ respectively. Moreover, it is also shown in [54, §3.2.1] that $\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}$ is a locally closed subset of $\mathscr{S}(\mathbf{G},X)_{K,\overline{\mathbb{F}}_{p}}^{\mathrm{tor}}$, and $\mathscr{C}_{\mathbb{X}}^{*}$ is a locally closed subset of $\mathscr{S}(\mathbf{G},X)_{K,\overline{\mathbb{F}}_{p}}^{*}$. Let $Z$ be a cusp label at level $K(N)$. This determines a locally closed boundary stratum $\mathscr{C}_{\mathbb{X},Z}\subset\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}$.
@@
The Igusa variety $\mathrm{Ig}^{b}$ over $\mathscr{C}_{\mathbb{X}}$ extends to a perfect scheme $\mathrm{Ig}^{b,\mathrm{tor}}$ over $\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}$. More precisely, we can define $\mathrm{Ig}^{b,\mathrm{tor}}$ as follows. Let $\mathcal{A}$ denote the universal semi-abelian scheme over $\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}$. This is the restriction to $\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}$ of the universal semi-abelian scheme over $\mathscr{S}(\mathbf{G},X)^{\mathrm{tor}}_{K,\overline{\mathbb{F}}_{p}}$. Then, we know from [54, Proposition 3.2.6] (which is exactly the same argument as in [8, Proposition 3.2.1]) that both the connected part $\mathcal{A}[p^{\infty}]^{\circ}$ of $\mathcal{A}[p^{\infty}]$ and the multiplicative part $\mathcal{A}[p^{\infty}]^{\mu}$ are $p$-divisible groups. We thus let $\mathcal{A}[p^{\infty}]^{(0,1)}=\mathcal{A}[p^{\infty}]^{\circ}/\mathcal{A}[p^{\infty}]^{\mu}$ be the biconnected part. We can similarly define $\mathbb{X}^{\circ},\mathbb{X}^{(0,1)}$ as the connected and biconnected parts of $\mathbb{X}$.
Following [8, Definition 3.2.9] we can define $\mathrm{Ig}^{b,\mathrm{tor}}$ to be the functor which, for a perfect $\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}$-scheme $\mathscr{T}$, parametrizes
  1. (1)
    $\mathcal{O}_{B_{\mathbb{Q}_{p}}}$ -linear isomorphisms $\rho:\mathcal{A}[p^{\infty}]^{\circ}\times_{\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}}\mathscr{T}\xrightarrow{\sim}\mathbb{X}^{\circ}\times_{\bar{\mathbb{F}}_{p}}{\mathscr{T}}$
  2. (2)
@@
    A scalar in $\mathbb{Z}^{\times}_{p}(\mathscr{T})$ such that the induced isomorphism $\rho^{(0,1)}:\mathcal{A}[p^{\infty}]^{(0,1)}\times_{\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}}\mathscr{T}\xrightarrow{\sim}\mathbb{X}_{\mathscr{T}}^{(0,1)}$ obtained by quotienting by the multiplicative parts commutes with the polarizations up to the given element of $\mathbb{Z}_{p}^{\times}(\mathscr{T})$.
Here, $\mathbb{Z}_{p}^{\times}(\mathscr{T})$ is the set of $\mathscr{T}$-points of the group scheme $\mathbb{Z}_{p}^{\times}\times\mathscr{C}_{\mathbb{X}}^{\mathrm{tor}}$. As explained in [8, §3.2.3-§3.2.7], this is representable by a perfect scheme.
@@
We define the partial minimal compactification $\mathrm{Ig}^{b,*}$ as the normalization of $\mathscr{C}_{\mathbb{X}}^{*}$ in $\mathrm{Ig}^{b}$. Since we have $\mathrm{Ig}^{b}\subset\mathrm{Ig}^{b,*},\mathrm{Ig}^{b,\mathrm{tor}}$, we will denote the boundaries by $\partial\mathrm{Ig}^{b,*}$ and $\partial\mathrm{Ig}^{b,\mathrm{tor}}$, respectively. These schemes are all perfect, and we can lift them to $p$-adic formal schemes $\mathfrak{Ig}^{b,*}$, $\mathfrak{Ig}^{b,\mathrm{tor}}$, $\partial\mathfrak{Ig}^{b,*}$, and $\partial\mathfrak{Ig}^{b,\mathrm{tor}}$ over $\mathrm{Spf}(W(\overline{\mathbb{F}}_{p}))$. We similarly denote by $\mathfrak{Ig}_{C}^{b,*}$, $\mathfrak{Ig}_{C}^{b,\mathrm{tor}}$, $\partial\mathfrak{Ig}_{C}^{b,*}$, and $\partial\mathfrak{Ig}_{C}^{b,\mathrm{tor}}$ the associated perfectoid spaces over $C$.